# ASE: Practical Acoustic Speed Estimation Beyond Doppler via Sound Diffusion Field

### v1

@@ 5.5 Case Study
Human Activity Recognition: Speed profiles act as the core of human activity recognition [48, 42, 52, 15, 79]. As shown in Fig. 26, the speed profiles (the dashed lines) generated by ASE can effectively distinguish different daily activities such as standing up, sitting down, and picking up objects. For instance, the speed profile of standing up shows a slower increase compared to sitting down. Picking up an object, however, forms a ’valley’ in the profile, reflecting the bending and rising motion. Fig. 26 provides an example of recognizing different activities through continuous monitoring. We keep it as a future direction to recognize different activities with ASE deployed on commodity smart speakers.
@@
Fall Detection: Fall detection is a vital yet challenging task. Passive speed estimation provides an effective way to detect falls [21, 91, 24, 33], mainly because falls induce a distinct speed pattern compared to normal daily activities. As shown in Fig. 26 and Fig. 26, a fall induces an unique pattern in the ACF matrix. The speed changes rapidly from a lower rate to a significantly higher rate within a brief period, yielding a large instantaneous acceleration. Following speed suggests further movement on the ground, before eventually decreasing to zero.
@@
Based on these observations, we employ deep learning for proof-of-concept fall detection. As the ACF matrix contains speed information, we leverage pretrained MobileNet-v2 [51] as the ACF matrix encoder to classify the fall events. To increase data availability, we develop a recording platform that connects multiple off-the-shelf audio devices to a server via MQTT [34]. In total, we have 305 samples, with 175 representing fall events. We achieve the F-1 score of 0.92, demonstrating the effectiveness of ASE for feature extractions.
@@
Gait Analysis: Gait analysis is important in healthcare, especially for patients with Parkinson’s disease [7, 81, 13]. Intuitively, the walking speeds estimated by ASE allow us to perform gait analysis. As a proof-of-concept, we recognize gait steps and cycles by detecting the peak in the speed profile, as shown in Fig. 26. Our analysis of 10 traces reports an average gait cycle time of 1.00s with a variance of 1.2ms, demonstrating the potential of ASE for gait applications.
6 Discussions
@@
Multiple Speeds: ASE assumes the human body shares the same speed $v$ for walk speed estimation, which is also adopted by many other works [89, 30, 41]. Eq. (12) suggests the potential to decompose the ACFs to account for varying speeds, which indicates an opportunity for future research to refine speed estimation by considering multiple speed scenarios.
@@
Multi-target Scenario: Current ASE is not designed for multi-person scenarios. In most cases, indoor walking speed downstream tasks, such as fall detection and gait monitoring, involve single person. Audio devices enable the interaction in real-time and alarm when abnormal conditions are detected. Therefore, ASE can already foster many applications. Yet it is also potential extend it to multi-person sensing in the future, e.g., by separating multipath signals for different ranges.
@@
Multiple Speakers: In ASE, we use OTDM with a single speaker and two channels, offering practical room-scale sensing. While this setup is common, OTDM can be extended to multiple speakers with advanced modulation. Future work can explore OTDM+OFDM design for higher CSI rates.
@@
Wi-Fi Speed Estimation: Numerous studies have utilized Wi-Fi for velocity capture [72, 41, 31]. However, the acoustic method offers several distinct advantages. Typically, commodity acoustic devices come with co-located microphone and speaker, allowing sensing with a single device, e.g., Amazon Echo Dot, without any hardware/firmware changes. In contrast, Wi-Fi sensing generally necessitates two separate transceivers and relies on specific chipsets (e.g., Intel 5300, Qualcomm Atheros NICs) that require firmware adjustments. In addition, acoustic devices bring side benefits by inherently making a voice interface available [78, 65, 92, 87]. For instance, besides running acoustic fall detection, a smart speaker can facilitate emergency calls directly. Therefore, it is imperative to enable accurate speed estimation for acoustic sensing. This work is orthogonal to existing Wi-Fi-based approaches.
@@
Phase-based Estimation: Recent studies [89, 31] have employed phase difference to address the challenges posed by high speed scenarios. They implicitly acquire speed by calculating acceleration or displacement. In contrast, our proposed framework, OTDM plus sound diffusion model, directly estimates speed without the need for further transformations. While our current approach utilizes channel amplitude, it is promising to extend this design to incorporate phase-based speed estimation, which we intend to explore in future work.
7 Related Works
@@
Acoustic Sensing: Acoustic sensing has been extensively explored recently and realize multiple applications, including indoor localization [32, 18, 36, 38], fine-grained gesture tracking [68, 40, 35, 83, 55, 60, 89, 30, 82], vital signs monitoring [66, 29, 45, 59, 61, 62], silent speech enhancement [54, 17], sound source localization [53, 64, 14], and communication [90, 11]. The latest work VeCare [88] introduces statistical acoustic sensing for presence detection, by extrapolating statistical WiFi sensing. ASE significantly differs from VeCare in multiple important aspects: 1) VeCare neither constructs the comprehensive model nor considers the fundamental differences between acoustic and electromagnetic waves. Conversely, ASE integrates a novel sound diffusion model with the acoustic channel for speed estimation for the first time. 2) VeCare mainly targets motion detection and breathing rate estimation, but does not address speed estimation (despite limited preliminary explorations). Differently, ASE not only establishes the theoretical model but also builds practical techniques for location-independent and large-coverage speed estimation, achieving remarkably better performance. 3) Additionally, we propose the first-of-its-kind OTDM design to increase the CSI rate, enabling previously challenging high speed estimation in a unique manner. Other than the key differences, we believe ASE complements VeCare towards a comprehensive framework for motion detection, vital sign monitoring, and speed estimation, in a distinct paradigm differing from the prevalent literature in acoustic sensing.
@@
Speed Estimation: Speed is vital information in human sensing, and has been exploited in various applications like fall detection [23, 43, 69, 33], interactive games [47], gait monitoring [58, 72, 81], and localization [46, 38]. Speed estimation, however, is a challenging and enduring task. Camera-based approaches, such as VICON [6], require professional devices and specialized calibration, which cannot be used in ubiquitous settings. In wireless sensing, speed is generally derived using DFS, regardless of Wi-Fi [46, 77, 56], acoustic [84, 35, 55, 89, 35, 53, 64], or mmWave radar [8, 74] signals. Recently, some work [89, 31] also acknowledge the problems of low CSI rate and have proposed using phase difference to estimate acceleration or displacement, indirectly inferring speed. However, they do not jump out of the DFS framework, which inherently limits observations to radial speed components. Moreover, inferring speed from displacement or acceleration introduces additional transformations, thereby increasing cumulative error [35]. In contrast, ASE introduces a novel and explicit speed estimation framework that statistically aggregates all paths by modeling the sound diffusion field, thereby overcoming these limitations. Despite the advances in statistical wireless sensing [86, 72], sound waves and EM waves differ in nature, and acoustic speed estimation incurs different challenges. We build the acoustic diffusion speed estimation model plus OTDM design, a new way of acoustic speed estimation.
8 Conclusions
@@
In this paper, we present ASE, a system for acoustic speed estimation. We propose a novel OTDM scheme to achieve a higher CSI rate for speed estimation. Inspired by the sound diffusion field, we establish a comprehensive theoretical model that enables full speed estimation from the spatial correlation of sounds. We extensively evaluate ASE on commodity devices, which achieves an average accuracy of 13 cm/s for normal walking speed estimation. Overall, the proposed ASE goes beyond the DFS-based paradigm and can open up new directions in acoustic sensing.

### v2

@@ 5.5. Case Study
Human Activity Recognition: Speed profiles act as the core of human activity recognition (Raeis et al., 2021; Ouyang et al., 2022; Shen et al., 2018; Ferlini et al., 2021; Xu et al., 2023). As shown in Fig. 29, the speed profiles (the dashed lines) generated by ASE can effectively distinguish different daily activities such as standing up, sitting down, and picking up objects. For instance, the speed profile of standing up shows a slower increase compared to sitting down. Picking up an object, however, forms a ’valley’ in the profile, reflecting the bending and rising motion. Fig. 29 provides an example of recognizing different activities through continuous monitoring. We keep it as a future direction to recognize different activities with ASE deployed on commodity smart speakers.
@@
Fall Detection: Fall detection is a vital yet challenging task. Passive speed estimation provides an effective way to detect falls (Hu et al., 2020a; Zheng et al., 2019; Ji et al., 2022; Lian et al., 2021), mainly because falls induce a distinct speed pattern compared to normal daily activities. As shown in Fig. 29 and Fig. 29, a fall induces a unique pattern in the ACF matrix. The speed changes rapidly from a lower rate to a significantly higher rate within a brief period, yielding a large instantaneous acceleration. Following speed suggests further movement on the ground, before eventually decreasing to zero.
@@
Based on these observations, we employ deep learning for proof-of-concept fall detection. As the ACF matrix contains speed information, we leverage pretrained MobileNet-v2 (Sandler et al., 2018) as the ACF matrix encoder to classify the fall events. To increase data availability, we develop a recording platform that connects multiple off-the-shelf audio devices to a server via MQTT (Light, 2017). In total, we have 305 samples, with 175 representing fall events. We achieve the F-1 score of 0.92, demonstrating the effectiveness of ASE for feature extractions.
@@
Gait Analysis: Gait analysis is important in healthcare, especially for patients with Parkinson’s disease (Akhtaruzzaman et al., 2016; Xu et al., 2019b; Chiang et al., 2020). Intuitively, the walking speeds estimated by ASE allow us to perform gait analysis. As a proof-of-concept, we recognize gait steps and cycles by detecting the peak in the speed profile, as shown in Fig. 29. Our analysis of 10 traces reports an average gait cycle time of 1.00s with a variance of 1.2ms, demonstrating the potential of ASE for gait applications.
6. Discussions
We conclude some discussions and limitations for further exploration.
@@ 6. Discussions
Multiple Speeds: ASE assumes the human body shares the same speed $v$ for walk speed estimation. The rationale behind our assumption is that we target room-scale sensing with relatively large coverage of several meters. In our proposed cases, the body speed dominates limb speeds. Moreover, a similar assumption is also implicitly adopted by many other works (Zhang et al., 2023b; Li et al., 2022c; Niu et al., 2022), i.e., hand-gesture works usually neglect body motions. As the next step, we are actively exploring multi-speed ASE. Potentially, Eq. (12) suggests the potential to decompose the ACFs to account for varying speeds, which indicates an opportunity for future research to refine speed estimation by considering multiple speed scenarios.
@@
Multi-target Scenario: Current ASE is not designed for multi-person scenarios. In most cases, indoor walking speed downstream tasks, such as fall detection and gait monitoring, involve a single person. Audio devices enable interaction in real-time and alert when abnormal conditions are detected. Therefore, ASE can already foster many applications. Yet it is also potentially extend it to multi-person sensing in the future, e.g., by separating multipath signals for different ranges. We also look forward to future work addressing the issue of multi-person movement.
@@
Multiple Speakers: In ASE, we use OTDM with a single speaker and two channels, offering practical room-scale sensing. While this setup is common, OTDM can be extended to multiple speakers with advanced modulation scheme. Future work can explore OTDM+OFDM design for higher CSI rates.
@@
Wi-Fi Speed Estimation: Numerous studies have utilized Wi-Fi for velocity capture (Wu et al., 2021; Niu et al., 2022; Li et al., 2024). However, the acoustic method offers several distinct advantages. Typically, commodity acoustic devices come with a co-located microphone and speaker, allowing sensing with a single device, e.g., Amazon Echo Dot, without any hardware/firmware changes. In contrast, Wi-Fi sensing generally necessitates two separate transceivers and relies on specific chipsets (e.g., Intel 5300, Qualcomm Atheros NICs) that require firmware adjustments. In addition, acoustic devices bring side benefits by inherently making a voice interface available (Xu et al., 2022; Wang et al., 2019a; Zhou et al., 2021; Zhang et al., 2016). For instance, besides running acoustic fall detection, a smart speaker can facilitate emergency calls directly. Therefore, it is imperative to enable accurate speed estimation for acoustic sensing. This work is orthogonal to existing Wi-Fi-based approaches.
@@
Phase-based Estimation: Recent studies (Zhang et al., 2023b; Li et al., 2024) have employed phase difference to address the challenges posed by high speed scenarios. They implicitly acquire speed by calculating acceleration or displacement. In contrast, our proposed framework, OTDM plus sound diffusion model, directly estimates speed without the need for further transformations. While our current approach utilizes channel amplitude, it is promising to extend this design to incorporate phase-based speed estimation, which we intend to explore in future work.
7. Related Works
@@
Acoustic Sensing: Acoustic sensing has been extensively explored recently and realize multiple applications, including indoor localization (Lian et al., 1 01; Gao et al., 2022; Mao et al., 2020, 2017), fine-grained gesture tracking (Wang et al., 2016; Nandakumar et al., 2016; Mao et al., 2016; Yun et al., 2015; Sun et al., 2018; Wang and Gollakota, 2019; Zhang et al., 2023b; Li et al., 2022c; Yang et al., 2023), vital signs monitoring (Wang et al., 2018; Li et al., 2022b; Qian et al., 2018a; Wan et al., 2021; Wang et al., 2019b, 2023; Kim et al., 2023), silent speech enhancement (Sun and Zhang, 2021; Fu et al., 2022; Yu et al., 2025), sound source localization (Shen et al., 2020; Wang and Sun, 2021; Fan et al., 2023; Veluri et al., 2024, 2023), and communication (Zhang et al., 2023c; Chen et al., 2022; Tian et al., 2025). The latest work VeCare (Zhang et al., 2023a) introduces statistical acoustic sensing for presence detection by extrapolating statistical WiFi sensing. ASE significantly differs from VeCare in multiple important aspects: 1) VeCare neither constructs the comprehensive model nor considers the fundamental differences between acoustic and electromagnetic waves. Conversely, ASE integrates a novel sound diffusion model with the acoustic channel for speed estimation for the first time. 2) VeCare mainly targets motion detection and breathing rate estimation, but does not address speed estimation (despite limited preliminary explorations). Differently, ASE not only establishes the theoretical model but also builds practical techniques for location-independent and large-coverage speed estimation, achieving remarkably better performance. 3) Additionally, we propose the first-of-its-kind OTDM design to increase the CSI rate, enabling previously challenging high speed estimation in a unique manner. Other than the key differences, we believe ASE complements VeCare towards a comprehensive framework for motion detection, vital sign monitoring, and speed estimation, in a distinct paradigm differing from the prevalent literature in acoustic sensing.
@@
Speed Estimation: Speed is vital information in human sensing, and has been exploited in various applications like fall detection (Hu et al., 2020b; Palipana et al., 2018; Wang et al., 2017; Lian et al., 2021), interactive games (Qian et al., 2017), gait monitoring (Umair Bin Altaf et al., 2015; Wu et al., 2021; Xu et al., 2019b), and localization (Qian et al., 2018b; Mao et al., 2017). Speed estimation, however, is a challenging and enduring task. Camera-based approaches, such as VICON (Vic, 2023), require professional devices and specialized calibration, which cannot be used in ubiquitous settings. In wireless sensing, speed is generally derived using DFS, regardless of Wi-Fi (Qian et al., 2018b; Xiong and Jamieson, 2013; Sun et al., 2015), acoustic (Yun et al., 2015; Mao et al., 2016; Sun et al., 2018; Zhang et al., 2023b; Mao et al., 2016; Shen et al., 2020; Wang and Sun, 2021), or mmWave radar (Blanco et al., 2022; Wu et al., 2020b) signals. Recently, some work (Zhang et al., 2023b; Li et al., 2024) also acknowledge the problems of low CSI rate and have proposed using phase difference to estimate acceleration or displacement, indirectly inferring speed. However, they do not jump out of the DFS framework, which inherently limits observations to radial speed components. Moreover, inferring speed from displacement or acceleration introduces additional transformations, thereby increasing cumulative error (Mao et al., 2016). In contrast, ASE introduces a novel and explicit speed estimation framework that statistically aggregates all paths by modeling the sound diffusion field, thereby overcoming these limitations. Despite the advances in statistical wireless sensing (Zhang et al., 2018; Wu et al., 2021), sound waves and EM waves differ in nature, and acoustic speed estimation incurs different challenges. We build the acoustic diffusion speed estimation model plus OTDM design, a new way of acoustic speed estimation.
8. Conclusions
@@
In this paper, we present ASE, an end-to-end system for acoustic speed estimation. We identify two research problems in the acoustic speed estimation: insufficient CSI rate to capture the speed and the oversimplified signal model that fails to capture the entire speed. To this end, we propose a novel OTDM scheme to achieve a higher CSI rate for speed estimation. Inspired by the sound diffusion field, we establish a comprehensive theoretical model that enables full speed estimation from the spatial correlation of sounds. We extensively evaluate ASE on commodity devices, which achieves an average accuracy of 13 cm/s for normal walking speed estimation. Overall, the proposed ASE goes beyond the DFS-based paradigm and can open up new directions in acoustic sensing.
Acknowledgements.
We sincerely thank all the anonymous reviewers for their valuable feedback throughout the submission process. We also wish to express our gratitude to the staff members at HKU who generously provided the initial space for our experiments. In addition, we are grateful to the lab members who assisted with proofreading the manuscript. This work is supported in part by NSFC under Grant No. 62222216, Hong Kong RGC GRF under Grant No. 17212224 and Healthy Longevity Catalyst Awards under Grant No. HLCA/E-712/22.